The density of 56% by weight 1_propanol is 0.8975 g/cm3.what is mole fraction of compound?

Grade:11

1 Answers

Ashutosh Sharma
8 years ago
mole fraction will be 0.276

consider a standered solution of 100gm as 1-propanol is 56% by weight and remaining 44% will be the water.
let,weight of 1-popanol be 56gm and its M.W. is 60.
so, number of moles will be= 56/60 =0.933
similarly, number of moles of water= 44/18= 2.44

so, the mole fracction of 1-propanol = 0933/(2.44+0.933)=0.276